OpenStudy (valirium):

2x + 5y= -4 4x - 5y = 22 What is the solution to the above system of equations?

OpenStudy (rcp031):

METHOD OF SUBSTITUTION AND ELIMINATION ! step 1 ELIMINATION of any variable x or y 2x + 5y= -4 *2 4x - 5y = 22 *1 4x+10y=-8 4x - 5y = 22 - + - reverse the operator symbol 0x+15y=-30 y=-30/15=-2 step 2 SUBSTITUTION 2x+5*-2=-4 2x-10=-4 2x=6 x=3. & y=-2

OpenStudy (lowkey.s):

@Valirium i can help